$SPAD/src/input ideal.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool ideal.output
)set message test on
)set message auto off
)clear all
--S 1 of 18
(n,m) : List DMP([x,y],FRAC INT)
--R
--R                                                                   Type: Void
--E 1
--S 2 of 18
m := [x**2+y**2-1]
--R
--R
--R          2    2
--R   (2)  [x  + y  - 1]
--R       Type: List(DistributedMultivariatePolynomial([x,y],Fraction(Integer)))
--E 2
--S 3 of 18
n := [x**2-y**2]
--R
--R
--R          2    2
--R   (3)  [x  - y ]
--R       Type: List(DistributedMultivariatePolynomial([x,y],Fraction(Integer)))
--E 3
--S 4 of 18
id := ideal m + ideal n
--R
--R
--R          2   1  2   1
--R   (4)  [x  - -,y  - -]
--R              2      2
--IType: PolynomialIdeals(Fraction(Integer),...
--E 4

--S 5 of 18
zeroDim? id
--R
--R
--R   (5)  true
--R                                                                Type: Boolean
--E 5
--S 6 of 18
zeroDim?(ideal m)
--R
--R
--R   (6)  false
--R                                                                Type: Boolean
--E 6
--S 7 of 18
dimension ideal m
--R
--R
--R   (7)  1
--R                                                        Type: PositiveInteger
--E 7
--S 8 of 18
(f,g):DMP([x,y],FRAC INT)
--R
--R                                                                   Type: Void
--E 8
--S 9 of 18
f := x**2-1
--R
--R
--R         2
--R   (9)  x  - 1
--R             Type: DistributedMultivariatePolynomial([x,y],Fraction(Integer))
--E 9
--S 10 of 18
g := x*(x**2-1)
--R
--R
--R          3
--R   (10)  x  - x
--R             Type: DistributedMultivariatePolynomial([x,y],Fraction(Integer))
--E 10
--S 11 of 18
relationsIdeal [f,g]

--R
--R
--R              2     3     2          2          3
--R   (11)  [- %B  + %A  + %A ] | [%A= x  - 1,%B= x  - x]
--IType: SuchThat(List(Polynomial(Fraction(Integer))),...
--E 11
--S 12 of 18
l: List DMP([x,y,z],FRAC INT)
--R
--R                                                                   Type: Void
--E 12
--S 13 of 18
l:=[x**2+2*y**2,x*z**2-y*z,z**2-4]
--R
--R
--R           2     2    2        2
--R   (13)  [x  + 2y ,x z  - y z,z  - 4]
--R     Type: List(DistributedMultivariatePolynomial([x,y,z],Fraction(Integer)))
--E 13
--S 14 of 18
ld:=primaryDecomp ideal l
--R
--R
--R               1    2             1    2
--R   (14)  [[x + - y,y ,z + 2],[x - - y,y ,z - 2]]
--R               2                  2
--IType: List(PolynomialIdeals(Fraction(Integer),...
--E 14
--S 15 of 18
reduce(intersect,ld)
--R
--R
--R              1      2  2
--R   (15)  [x - - y z,y ,z  - 4]
--R              4
--IType: PolynomialIdeals(Fraction(Integer),...
--E 15
--S 16 of 18
reduce(intersect,[radical ld.i for i in 1..2])
--R
--R
--R               2
--R   (16)  [x,y,z  - 4]
--IType: PolynomialIdeals(Fraction(Integer),...
--E 16

--S 17 of 18
radical ideal l
--R
--R
--R               2
--R   (17)  [x,y,z  - 4]
--IType: PolynomialIdeals(Fraction(Integer),...
--E 17
--S 18 of 18
quotient(ideal l,y)
--R
--R
--R               2
--R   (18)  [x,y,z  - 4]
--IIype: PolynomialIdeals(Fraction(Integer),...
--E 18
)spool
)lisp (bye)

———-

References^{[1] nothing}

7